OpenStudy (anonymous):

What is the molarity of a saturated solution of potassium sulfate if the solubility is 13 g per 100 g H2O at 25 °C? The density of the solution is 1.1g/mL

OpenStudy (souvik):

what is the chemical formula of potassium sulfate?

OpenStudy (souvik):

Molar mass of potassium sulfate: 174.259 g/mol the solubility of \(K _2SO _4\) is13 g per 100 g at room temperature... so total mass of the solution=100+13=113 g the volume of the solution=113/1.1=102.727 mL so there is 13 g \(K _2SO _4\) in 102.727 mL of the solution or there is 13/174.259=0.075 mole of \(K _2SO _4\) so the molarity is 0.075/102.727*1000=0.726(M)